OpenStudy (anonymous):

Evaluate 10 + 8 ÷ 2 – 4.

OpenStudy (anonymous):

10+4-4= 14-4 = 10

OpenStudy (muscrat123):

yea okay

OpenStudy (muscrat123):

?

OpenStudy (muscrat123):

thats not how u solve

OpenStudy (muscrat123):

you should right it out. yes it is correct, but i recommend writing out the steps to make less mistakes and u r not supposed to give direct answers on OS a